Estimate the vapor pressure of carbon dioxide at 2
Answer #1

a) the equation of redich kwong is RT so we have than

2 2.5 9(V2P .42748 Pe

RTC = 0.08664 3 P

now than we have the equations need the value so

R = 0.082atm l /mol K for the CO2 the pressure critc is 72.8atm and your Tc is 30.98C = 303.98K

Vm = RT to 284K the Vm = 0.082l/mol K x 284IK = 23.288l

a = \frac{1}{9(\sqrt[3]{2}-1)}\frac{0.082^{2}303.98^{2.5}}{72.81} = 042748\frac{0.082^{2}303.98^{2.5}}{72.81}

a = 63.60

b = 0.08664\frac{0.082x303.98}{72.81}

b = 0.030

RT remember than T is T ambien in this case 284K

only replacing value and get

P = 7.33x10-3atm

the rault laws say than the P partial of a gas = Pure for your molar fraction so

Pco = P Xco as the cocacola is a 4% M of CO2 yout X = 4/100 = 0.04 so your Partial pressure is

Pco = 0.04x7.33x10-3atm = 3x10-4atm